OpenStudy (nincompoop):

find the area between the cuber and the x-axis from x=-2 to x=1 of definite integral f(x)= x^3 + x^2

OpenStudy (nincompoop):

\[\int\limits_{-2}^{1}(x^3+x^2)dx=\int\limits_{-2}^{1}x^3dx+\int\limits_{-2}^{1}x^2dx\]

OpenStudy (nincompoop):

\[=\frac{ x^4 }{4 }+\frac{ x^3 }{ 3 }\]